janet is three years older than her sister julie janets brother is eight years younger than their sister julie the sum of all of their ages is 55. How do I write this in symbolic language algebra?
Answer by ankor@dixie-net.com(22740)

let x = Janet's age
let y = Julie's age
let z = brother's age
:
Punctuation is a wonderful thing! Write an equation for each statement.
:
Janet is three years older than her sister, Julie.
x = y + 3
Janet's brother is eight years younger than their sister, Julie.
z = y - 8
The sum of all of their ages is 55.
x + y + z = 55